OpenStudy (anonymous):

find the exact value sin^-1[cos2pi/3]

OpenStudy (anonymous):

sin^-1(-1/2)=210 330=11pi/6 right

OpenStudy (anonymous):

i dont got it

myininaya (myininaya):

your number is outside the range

myininaya (myininaya):

\[\frac{11\pi}{6}-2\pi \]

myininaya (myininaya):

which you can simplify to an element in the range

myininaya (myininaya):



myininaya (myininaya):